Exercise 15-4 (Algo) Sales-type lease; lessor; balance sheet and income statement effects [LO15-2]

On June 30, 2021, Georgia-Atlantic, Inc. leased warehouse equipment from IC Leasing Corporation. The lease agreement calls for Georgia-Atlantic to make semiannual lease payments of $403,067 over a five-year lease term (also the asset’s useful life), payable each June 30 and December 31, with the first payment at June 30, 2021. Georgia-Atlantic's incremental borrowing rate is 8%, the same rate IC used to calculate lease payment amounts. IC purchased the equipment from Builders, Inc. at a cost of $3.4 million. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. What pretax amount related to the lease would IC report in its balance sheet at December 31, 2021?
2. What pretax amount related to the lease would IC report in its income statement for the year ended December 31, 2021?
(For all requirements, enter your answers in whole dollars and not in millions. Round your final answers to the nearest whole dollar.)
Answer #1

Solution 1:

Semiannual lease payment = $403,067

Total semiannual payments = 5*2 = 10

Incremental borrowing rate = 8%, 4% semiannual

Present value of minimum lease payments = Semi Annual lease payments * Cumulative PV Factor of annuity due for 10 periods at 4%

= $403,067 * 8.435332 = $3,400,003

Semiannual payment on 30.06.2021 = $403,067

Amount receivable on 30.06.2021 = ($3,400,003 - $403,067) = $2,996,936

Interest revenue for 31.12.2021 = $2,996,936 * 4% = $119,877

Semiannual lease payment on 31.12.2021 = $403,067

Pre tax amount of net receivables IC report in its balance sheet at December 31, 2021 = $2,996,936 + $119,877 - $403,067 = $2,713,746

Solution 2:

Pre tax amount of interest revenue IC report in its income statement for year ended Dec 31, 2021 = $2,996,936 * 4% = $119,877